# Time Evolution and Thermal Renormalization Group Flow in Cosmology

### v2

@@ Acknowledgements.
with $\lambda_{k}=\Lambda_{k}k^{-2}$, $g_{k}=Gk^{2}$ is the dimensionful Newton constant, and $G$ is scale-dependent due to the absence of quantum gravity, so $k\partial_{k}g_{k}=2g_{k}$. By using dimensionless couplings, $g_{0,k}=\tilde{g}_{0,k}k^{4}$ and $g_{2,k}=\tilde{g}_{2,k}k^{2}$ the flow equation for the field independent term is substituted into (16) and the thermal RG flow equation for the cosmological constant is obtained.
As a final step, let us rewrite the flow equations for dimensionless couplings $g_{0,k}=\tilde{g}_{0,k}k^{4}$, $g_{2,k}=\tilde{g}_{2,k}k^{2}$ and $g_{4,k}=\tilde{g}_{4,k}$ by using the subtracted form (14). In this case one finds,
$(4+k\partial_{k})$ $\tilde{g}_{0,k}=\frac{1}{6\pi^{2}}\frac{\coth{\left(\frac{\sqrt{1+\tilde{g}_{2,k}}}{2\tau}\right)}}{2\sqrt{1+\tilde{g}_{2,k}}}-\frac{1}{12\pi^{2}}\coth{\frac{1}{2\tau}}-\frac{1}{24\pi^{2}}\frac{1+\tau\sinh{\frac{1}{\tau}}}{\tau-\tau\cosh{\frac{1}{\tau}}}\tilde{g}_{2,k},$ (17)
$(2+k\partial_{k})$ $\tilde{g}_{2,k}=\frac{1}{6\pi^{2}}\left[-\frac{\tilde{g}_{4,k}\coth\left(\frac{\sqrt{1+\tilde{g}_{2,k}}}{2\tau}\right)}{4\left(1+\tilde{g}_{2,k}\right)^{3/2}}-\frac{\tilde{g}_{4,k}\;\text{csch}^{2}\left(\frac{\sqrt{1+\tilde{g}_{2,k}}}{2\tau}\right)}{8\tau\left(1+\tilde{g}_{2,k}\right)}\right],$ (18)
$k\partial_{k}$ $\tilde{g}_{4,k}=\frac{1}{6\pi^{2}}\left[\frac{9\tilde{g}_{4,k}^{2}\coth\left(\frac{\sqrt{1+\tilde{g}_{2,k}}}{2\tau}\right)}{8(1+\tilde{g}_{2,k})^{5/2}}+\frac{9\tilde{g}_{4,k}^{2}\text{csch}^{2}\left(\frac{\sqrt{1+\tilde{g}_{2,k}}}{2\tau}\right)}{16\tau(1+\tilde{g}_{2,k})^{2}}+\frac{3\tilde{g}_{4,k}^{2}\coth\left(\frac{\sqrt{1+\tilde{g}_{2,k}}}{2\tau}\right)\text{csch}^{2}\left(\frac{\sqrt{1+\tilde{g}_{2,k}}}{2\tau}\right)}{16\tau^{2}(1+\tilde{g}_{2,k})^{3/2}}\right].$ (19)

### v3

@@ Acknowledgements
with $\lambda_{k}=\Lambda_{{\mathrm{cosmo}},k}k^{-2}$, $g_{k}=Gk^{2}$ is the dimensionful Newton constant, and $G$ is scale-dependent due to the absence of quantum gravity, so $k\partial_{k}g_{k}=2g_{k}$. By using dimensionless couplings, $g_{0,k}=\tilde{g}_{0,k}k^{4}$ and $g_{2,k}=\tilde{g}_{2,k}k^{2}$ the flow equation for the field independent term is substituted into (16) and the thermal RG flow equation for the cosmological constant is obtained.
As a final step, let us rewrite the flow equations for dimensionless couplings $g_{0,k}=\tilde{g}_{0,k}k^{4}$, $g_{2,k}=\tilde{g}_{2,k}k^{2}$ and $g_{4,k}=\tilde{g}_{4,k}$ by using the subtracted form (A). In this case one finds,
$(4+k\partial_{k})\tilde{g}_{0,k}=$ $\frac{1}{6\pi^{2}}\frac{\coth{\left(\frac{\sqrt{1+\tilde{g}_{2,k}}}{2\tau}\right)}}{2\sqrt{1+\tilde{g}_{2,k}}}-\frac{1}{12\pi^{2}}\coth{\frac{1}{2\tau}}$
$-\frac{1}{24\pi^{2}}\frac{1+\tau\sinh{\frac{1}{\tau}}}{\tau-\tau\cosh{\frac{1}{\tau}}}\tilde{g}_{2,k},$ (17)
$(2+k\partial_{k})\tilde{g}_{2,k}=$ $\frac{1}{6\pi^{2}}\left[-\frac{\tilde{g}_{4,k}\coth\left(\frac{\sqrt{1+\tilde{g}_{2,k}}}{2\tau}\right)}{4\left(1+\tilde{g}_{2,k}\right)^{3/2}}\right.$ (18)
$\left.-\frac{\tilde{g}_{4,k}\;\text{csch}^{2}\left(\frac{\sqrt{1+\tilde{g}_{2,k}}}{2\tau}\right)}{8\tau\left(1+\tilde{g}_{2,k}\right)}\right],$
$k\partial_{k}\tilde{g}_{4,k}=$ $\frac{1}{6\pi^{2}}\left[\frac{9\tilde{g}_{4,k}^{2}\coth\left(\frac{\sqrt{1+\tilde{g}_{2,k}}}{2\tau}\right)}{8(1+\tilde{g}_{2,k})^{5/2}}\right.$ (19)
$\left.+\frac{9\tilde{g}_{4,k}^{2}\text{csch}^{2}\left(\frac{\sqrt{1+\tilde{g}_{2,k}}}{2\tau}\right)}{16\tau(1+\tilde{g}_{2,k})^{2}}\right.$
$\left.+\frac{3\tilde{g}_{4,k}^{2}\coth\left(\frac{\sqrt{1+\tilde{g}_{2,k}}}{2\tau}\right)\text{csch}^{2}\left(\frac{\sqrt{1+\tilde{g}_{2,k}}}{2\tau}\right)}{16\tau^{2}(1+\tilde{g}_{2,k})^{3/2}}\right].$